# Localized Diffusion Models for High Dimensional Distributions Generation

### v1

@@ Abstract
Diffusion models are the state-of-the-art tools for various generative tasks. However, estimating high-dimensional score functions makes them potentially suffer from the curse of dimensionality (CoD). This underscores the importance of better understanding and exploiting low-dimensional structure in the target distribution. In this work, we consider locality structure, which describes sparse dependencies between model components. Under locality structure, the score function is effectively low-dimensional, so that it can be estimated by a localized neural network with significantly reduced sample complexity. This motivates the localized diffusion model, where a localized score matching loss is used to train the score function within a localized hypothesis space. We prove that such localization enables diffusion models to circumvent CoD, at the price of additional localization error. Under realistic sample size scaling, we show both theoretically and numerically that a moderate localization radius can balance the statistical and localization error, leading to a better overall performance. The localized structure also facilitates parallel training of diffusion models, making it potentially more efficient for large-scale applications.
@@ 1 Introduction
Over the past decade, numerous neural network (NN)-based sampling algorithms have emerged in the machine learning literature, demonstrating remarkable performance across various tasks. These methods, often referred to as generative models, include approaches such as normalizing flows [35], variational encoders [24], generative adversarial network [18], and diffusion models [37, 21, 38]. Among all generative models, diffusion model (DM), usually referring to the denoising diffusion probabilistic model (DDPM) [21], is a state-of-the-art and widely used approach. It has gained great popularity due to its capability in generating high quality samples, particularly in tasks such as image synthesis [21, 14]. A series of recent studies [28, 9, 4, 32, 7, 43] theoretically justify the approximation and generalization capabilities of DMs over a broad class of target distributions. However, there remains limited understanding of its effectiveness in handling high-dimensional distributions.
@@
DMs are known to be expensive when it comes to training with high-dimensional data. The training sample size needs to grow exponentially with the problem dimension [39, 32], known as the curse of dimensionality (CoD) in the literature. Various attempts have been made to avoid it by leveraging low-dimensional structures within the target distribution. The manifold hypothesis [15] which assumes that the data lies on a low-dimensional manifold, is often envoked to postulate such structures. For such data, [32, 7, 40, 1] show that the sample complexity of DMs depends on the dimension of the manifold rather than the ambient dimension, and DMs can avoid the CoD with appropriate NN structure. There are also studies considering Gaussian mixtures [36, 45, 17] to avoid the CoD. In both manifold hypothesis and the Gaussian mixture models, although the ambient space is high dimension, there is a low-dimensional latent structure that effectively characterizes the target distribution.
@@
While these concepts of low effective dimension can cover many applications, there are still important cases left open. One large class of high-dimensional distributions are those with locality structure [5, 41, 16, 13]. We say a distribution has locality structure if each model component only has strong conditional dependence on a sparse selection of the other model components. An illustrative example is the Ginzburg-Landau model from statistical physics [27], where $d$ particles in one-dimensional configurations follow the distribution
$p(x_{1},\dots,x_{d})=\frac{1}{Z}\exp\left(\sum_{j=1}^{d}V(x_{j})+\sum_{j=1}^{d-1}W(x_{j},x_{j+1})\right).$
@@
In this model, each particle, denoted by $x_{j}$, interacts directly only with its neighbors $x_{j\pm 1}$. Such sparse dependence structure arises naturally in spatial models, and has been successfully applied in various fields such as spatial statistics [5], data assimilation [34], quantum mechanics [25] and sampling [41]. We refer to [13] for a detailed review on the locality structure.
@@
An important property of distributions with locality structure, or localized distributions, is that their score functions are effectively low-dimensional [41, 13]. Due to the conditional independence, the score component $s_{j}(x)=\nabla_{j}\log p(x)$ depends only on $x_{\mathcal{N}_{j}}$, where $x_{\mathcal{N}_{j}}$ is the component conditionally dependent on $x_{j}$. If the conditional dependencies are sparse, the dimension of $x_{\mathcal{N}_{j}}$ is much smaller than the ambient dimension $d$, so that the score function can be regarded as a collection of low-dimensional functions $\{s_{j}\}_{j\in[d]}$. This suggests that learning the score functions of localized distributions does not suffer from the CoD.
@@
Motivated by this, we propose the localized diffusion model (LDM), which embeds the locality structure into the hypothesis space of the score function, reducing a high dimensional score matching problem to a low dimensional one. With small effective dimension, the statistical error of score estimation is significantly reduced. On the other hand, localizing the hypothesis space introduces additional localization error. By a complete approximation and generalization analysis, we show that by adjusting the localization radius, one can balance the tradeoff between the statistical error and the localization error to achieve smaller overall error. This can be interpreted as a tradeoff between variance and bias. Such tradeoff is validated by numerical experiments on high-dimensional time series data. Finally, we find that LDM can be interpreted as a collection of diffusion models on low-dimensional marginals. That is, we construct the samplers by combining local samplers for the marginals of the localized distributions. This allows LDM to be trained parallelly, which is practically important for large-scale applications.
@@ 1.1 Related Work
Analysis of Diffusion Models
@@
Localized Sampler
@@ 1.2 Notations
  • Sets. Denote $[n]=\{1,2,\dots,n\}$, and the cardinality of a set $A$ as $|A|$. Given $x\in\mathbb{R}^{n}$ and $A\subset[n]$, denote $x_{A}$ as the subvector of $x$ with components’ indices from $A$.
@@ 2.1 Diffusion Models
    $\widehat{s}=\operatorname*{arg\,min}_{s_{\theta}}\mathcal{L}(s_{\theta}),$
@@
$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon_{t},\quad\epsilon_{t}\sim\mathsf{N}(0,I).$
@@ 2.2.2 Localized Distributions
$s_{j}(x)\approx\widehat{s}_{\theta,j}(x_{\mathcal{N}_{j}^{r}}).$
@@ 2.3 Locality Structure in Diffusion Models
$\nabla_{ij}^{2}\log p_{t}(x_{t})=\alpha_{t}^{2}\sigma_{t}^{-4}\mathsf{Cov}_{p_{0|t}(x_{0}|x_{t})}\left(x_{0,i},x_{0,j}\right).$
@@
(1) While 2.1 assumes log-concavity to apply 2.2, the exponential decay of correlations is ubiquitous and does not inherently depend on log-concavity. The assumption is adopted here for simplicity and to derive an explicit quantitative bound.
(2) It is natural to ask if 2.1 can be extended to the case where $p_{0}$ is only approximately localized. The answer depends on the sparsity of the graph $G$ and the decay rate of $\nabla^{2}_{ij}\log p_{0}$. The resulting bound will be complicated, and we do not pursue it here.
@@
$p(x)\propto\exp\left(\frac{1}{2}x^{\rm T}Ax-\frac{\gamma}{2}\left\|x\right\|^{2}\right),$
@@
    $-\nabla^{2}\log p(x)=-A+\gamma I=\frac{1}{h^{2}}\begin{bmatrix}2&-1&0&\cdots&0\\ -1&2&-1&\cdots&0\\ 0&-1&2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&2\end{bmatrix}+\gamma I.$
@@
    $\kappa=\frac{\gamma+4h^{-2}\sin^{2}\frac{d\pi}{2(d+1)}}{\gamma+4h^{-2}\sin^{2}\frac{\pi}{2(d+1)}}\approx\frac{\sin^{2}\frac{d\pi}{2(d+1)}}{\sin^{2}\frac{\pi}{2(d+1)}}\asymp d^{2}.$
  2. 2.
    Extended-domain (locality) type. Fix the mesh size $h=h_{0}$, and consider an extended domain $[0,(d+1)/h_{0}]$. Take $x_{k}=kh_{0}$, then $-\nabla^{2}\log p(x)$ has the same form as above with $h=h_{0}$. Therefore,
    $\kappa=\frac{\gamma+4h_{0}^{-2}\sin^{2}\frac{d\pi}{2(d+1)}}{\gamma+4h_{0}^{-2}\sin^{2}\frac{\pi}{2(d+1)}}\approx\frac{\gamma+4h_{0}^{-2}}{\gamma}\asymp 1.$
@@ 3.1.3 Localized Score Matching
^$\widehat{\mathcal{L}}_{N}(s_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\int_{\underline{t}}^{T}\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|s_{\theta}(\alpha_{t}X^{(i)}+\sigma_{t}\epsilon_{t},t)+\sigma_{t}^{-1}\epsilon_{t}\right\|^{2}\right]\mathrm{d}t.$ (3.6)
Notice ^$\widehat{\mathcal{L}}_{N}$ is decomposable: ^$\widehat{\mathcal{L}}_{N}(s_{\theta})=\sum_{j=1}^{b}\widehat{\mathcal{L}}_{j,N}(u_{\theta,j})$, where
^$\widehat{\mathcal{L}}_{j,N}(u_{\theta,j})=\frac{1}{N}\sum_{i=1}^{N}\int_{\underline{t}}^{T}\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|u_{\theta,j}(\alpha_{t}X_{\mathcal{N}_{j}^{r}}^{(i)}+\sigma_{t}\epsilon_{t,\mathcal{N}_{j}^{r}},t)+\sigma_{t}^{-1}\epsilon_{t,j}\right\|^{2}\right]\mathrm{d}t.$ (3.7)
@@ 3.2.1 Error Decomposition
$\mathbb{E}_{\mathsf{Q}}\left[\exp\left(\frac{1}{2}\int_{0}^{T-{\underline{t}}}\left\|\widehat{s}(Y_{t},T-t)-s(Y_{t},T-t)\right\|^{2}\mathrm{d}t\right)\right]<\infty,$
@@ 3.2.2 Localized Score Function
$s^{*}=\operatorname*{arg\,min}_{s_{\theta}\in\mathscr{H}_{r}^{*}}\int_{\underline{t}}^{T}\int\left\|s_{\theta}(x,t)-s(x,t)\right\|^{2}p_{t}(x)\mathrm{d}x\mathrm{d}t$
$\Leftrightarrow~{}\forall j,~{}s^{*}_{j}(x,t)=u_{j}^{*}(x_{\mathcal{N}_{j}^{r}},t),\quad u_{j}^{*}=\operatorname*{arg\,min}_{u_{\theta,j}\in\mathscr{U}_{j}}\int_{\underline{t}}^{T}\int\|u_{\theta,j}(x_{\mathcal{N}_{j}^{r}},t)-s_{j}(x,t)\|^{2}p_{t}(x)\mathrm{d}x\mathrm{d}t.$
@@
$|\mathcal{N}_{j}^{r}|=|\{i\in\mathbb{Z}^{\nu}:\left\|i\right\|_{1}\leq r\}|\leq(2r+1)^{\nu}<1+(3r)^{\nu}.$
@@
$C=2S\max\{1,m^{-1}\}\nu!\kappa^{2\nu+1}\log\kappa,\quad c=-2\log(1-\kappa^{-1}).$
@@
$\int_{0}^{T}\left\|s_{\theta}(x,t)-s(x,t)\right\|_{L^{2}(p_{t})}^{2}\mathrm{d}t\leq Cd(r+1)^{\nu}\mathrm{e}^{-c(r+1)},$
@@ 3.3.1 Equivalent to Diffusion Models for Marginals
$\mathcal{L}_{j}(u_{\theta,j})=~{}$ $\mathbb{E}_{x_{0,\mathcal{N}_{j}^{r}}\sim p_{0}}\int_{\underline{t}}^{T}\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|u_{\theta,j}(\alpha_{t}x_{0,\mathcal{N}_{j}^{r}}+\sigma_{t}\epsilon_{t,\mathcal{N}_{j}^{r}},t)+\sigma_{t}^{-1}\epsilon_{t,j}\right\|^{2}\right]\mathrm{d}t$
[…]

### v2

@@ Abstract
Diffusion models are state-of-the-art tools for various generative tasks. Yet training these models involves estimating high-dimensional score functions, which in principle suffers from the curse of dimensionality. It is therefore important to understand how low-dimensional structure in the target distribution can be exploited in these models. Here we consider locality structure, which describes certain sparse conditional dependencies among the target random variables. Given some locality structure, the score function is effectively low-dimensional, so that it can be estimated by a localized neural network with significantly reduced sample complexity. This observation motivates the localized diffusion model, where a localized score matching loss is used to train the score function within a localized hypothesis space. We prove that such localization enables diffusion models to circumvent the curse of dimensionality, at the price of additional localization error. Under realistic sample size scaling, we then show both theoretically and numerically that a moderate localization radius can balance the statistical and localization errors, yielding better overall performance. Localized structure also facilitates parallel training, making localized diffusion models potentially more efficient for large-scale applications.
@@ 1 Introduction
Over the past decade, numerous neural network (NN)-based sampling algorithms have emerged in the machine learning literature, demonstrating remarkable performance in varied applications. These methods, often referred to as generative models, include normalizing flows [36], variational auto-encoders [24], generative adversarial networks [18], and diffusion models [38, 21, 39]. Among generative models, diffusion models (DMs)—a term usually referring to “denoising diffusion probabilistic models” (DDPMs) [21]—are a state-of-the-art and widely used approach. They have gained great popularity due to their ability to generate high-quality samples, particularly in tasks such as image synthesis [21, 14]. A series of recent studies [28, 9, 4, 33, 7, 44] theoretically justify the approximation and generalization capabilities of DMs over a broad class of target distributions. Understanding of their effectiveness in representing high-dimensional distributions, however, remains limited.
@@
DMs are known to be expensive when training with high-dimensional data. The training sample size needs to grow exponentially with the problem dimension [40, 33], a scaling known as the curse of dimensionality (CoD) in the literature. Various attempts have been made to avoid this scaling by leveraging low-dimensional structure within the target distribution. The manifold hypothesis [15], which postulates that the data lie on a low-dimensional manifold, is often invoked in these settings. For such data, [33, 7, 41, 1] show that the sample complexity of DMs depends on the dimension of the manifold rather than the ambient dimension, and that DMs can thus avoid the CoD with appropriate NN structure. There are also studies considering Gaussian mixtures [37, 46, 17] to avoid the CoD. In both manifold hypothesis and the Gaussian mixture models, although the ambient space is high-dimensional, there is a low-dimensional latent structure that effectively characterizes the target distribution.
@@
While these concepts of low “effective dimension” cover many applications, there are still important cases left open. One large class of high-dimensional distributions are those with locality structure [5, 42, 16, 13]. We say a joint distribution has locality structure if each coordinate random variable only has strong conditional dependence on a small subset of the other coordinates. An illustrative example is the Ginzburg–Landau model from statistical physics [27], where $d$ particles in one-dimensional configurations follow the distribution
$p(x_{1},\dots,x_{d})=\frac{1}{Z}\exp\left(\sum_{j=1}^{d}V(x_{j})+\sum_{j=1}^{d-1}W(x_{j},x_{j+1})\right).$ (1.1)
@@
In this model, each particle, denoted by $x_{j}$, interacts directly only with its neighbors $x_{j\pm 1}$. Such sparse dependence structure arises naturally in spatial models, and has been successfully applied in various fields such as spatial statistics [5], data assimilation [35], quantum mechanics [25], and sampling [42]. We refer to [13] for a detailed review on locality structure.
@@
An important property of distributions with locality structure, or localized distributions, is that their score functions are effectively low-dimensional [42, 13]. Due to conditional independence, the $j$th component of the score function, $s_{j}(x)=\nabla_{j}\log p(x)$, depends only on $x_{\mathcal{N}_{j}}$, where $x_{\mathcal{N}_{j}}$ is the component conditionally dependent on $x_{j}$. If the conditional dependencies are sparse, the dimension of $x_{\mathcal{N}_{j}}$ is much smaller than the ambient dimension $d$, so that the score function can be regarded as a collection of low-dimensional functions $\{s_{j}\}_{j\in[d]}$. This suggests that learning the score functions of localized distributions does not suffer from the CoD.
@@
Motivated by this, we propose the localized diffusion model (LDM), which embeds the locality structure into the hypothesis space of the score function, reducing a high-dimensional score matching problem to a low-dimensional one. With small effective dimension, the statistical error of score estimation is significantly reduced. On the other hand, localizing the hypothesis space introduces additional localization error. By a complete approximation and generalization analysis, we show that by adjusting the localization radius, one can balance the tradeoff between the statistical error and the localization error to achieve smaller overall error. This can be interpreted as a tradeoff between variance and bias. Such tradeoff is validated by numerical experiments on high-dimensional time series data. Finally, we find that LDM can be interpreted as a collection of diffusion models on low-dimensional marginals. That is, we construct the samplers by combining local samplers for the marginals of the localized distributions. This allows LDM to be trained in parallel, which is practically important for large-scale applications.
@@ 1.1 Related Work
1.1.1 Analysis of Diffusion Models
@@
Locality in DMs has also been studied from a different perspective, i.e., the mechanism of creativity or generalization capability in image generation task. [23, 32] suggest that the inductive bias introduced by localized denoisers in neural architectures such as convolutional neural networks may be the key for DMs to generalize rather than memorize. Although both works derive an empirical localized denoiser, it is introduced to explain the emergence of creativity, rather than to be used to improve sample complexity. Meanwhile, they focus on empirical studies rather than rigorous numerical and statistical analysis.
1.1.2 Localized Sampler
@@ 1.2 Notations
  • Sets. Denote $n]=\{1,2,\dots,n\}$, and the cardinality of a set $A$ as $|A|$. Given $x\in\mathbb{R}^{n}$ and $A\subset[n]$, denote $x_{A}$ as the subvector of $x$ with components’ indices from $A$.
@@ 2.1 Diffusion Models
    $\widehat{s}=\operatorname*{arg\,min}_{s_{\theta}}\mathcal{L}(s_{\theta}),$ (2.3)
@@
$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon_{t},\quad\epsilon_{t}\sim\mathsf{N}(0,I).$ (2.7)
@@ 2.2.2 Localized Distributions
$s_{j}(x)\approx\widehat{s}_{\theta,j}(x_{\mathcal{N}_{j}^{r}}).$ (2.18)
@@ 2.3 Locality Structure in Diffusion Models
$\nabla_{ij}^{2}\log p_{t}(x_{t})=\alpha_{t}^{2}\sigma_{t}^{-4}\mathsf{Cov}_{p_{0|t}(x_{0}|x_{t})}\left(x_{0,i},x_{0,j}\right).$ (2.21)
@@
While 2.1 assumes log-concavity to apply 2.2, the exponential decay of correlations is ubiquitous and does not inherently depend on log-concavity. The assumption is adopted here for simplicity and to derive an explicit quantitative bound.
@@
$p(x)\propto\exp\left(\frac{1}{2}x^{\rm T}Ax-\frac{\gamma}{2}\left\|x\right\|^{2}\right),$ (2.23)
@@
    $\nabla^{2}\log p(x)=-A+\gamma I=\frac{1}{h^{2}}\begin{bmatrix}2&-1&0&\cdots&0\\ -1&2&-1&\cdots&0\\ 0&-1&2&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&2\end{bmatrix}+\gamma I.$ (2.24)
@@
    $\kappa=\frac{\gamma+4h^{-2}\sin^{2}\frac{d\pi}{2(d+1)}}{\gamma+4h^{-2}\sin^{2}\frac{\pi}{2(d+1)}}\approx\frac{\sin^{2}\frac{d\pi}{2(d+1)}}{\sin^{2}\frac{\pi}{2(d+1)}}\asymp d^{2}.$ (2.25)
  • Extended-domain (locality) type. Fix the mesh size $h=h_{0}$, and consider an extended domain $0,(d+1)/h_{0}]$. Take $x_{k}=kh_{0}$, then $\nabla^{2}\log p(x)$ has the same form as above with $h=h_{0}$. Therefore,
    $\kappa=\frac{\gamma+4h_{0}^{-2}\sin^{2}\frac{d\pi}{2(d+1)}}{\gamma+4h_{0}^{-2}\sin^{2}\frac{\pi}{2(d+1)}}\approx\frac{\gamma+4h_{0}^{-2}}{\gamma}\asymp 1.$ (2.26)
@@ 3.1.3 Localized Score Matching
$\widehat{\mathcal{L}}_{N}(s_{\theta})=\frac{1}{N}\sum_{i=1}^{N}\int_{\underline{t}}^{T}\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|s_{\theta}(\alpha_{t}X^{(i)}+\sigma_{t}\epsilon_{t},t)+\sigma_{t}^{-1}\epsilon_{t}\right\|^{2}\right]\mathrm{d}t.$ (3.6)
Notice $\widehat{\mathcal{L}}_{N}$ is decomposable: $\widehat{\mathcal{L}}_{N}(s_{\theta})=\sum_{j=1}^{b}\widehat{\mathcal{L}}_{j,N}(u_{\theta,j})$, where
$\widehat{\mathcal{L}}_{j,N}(u_{\theta,j})=\frac{1}{N}\sum_{i=1}^{N}\int_{\underline{t}}^{T}\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|u_{\theta,j}(\alpha_{t}X_{\mathcal{N}_{j}^{r}}^{(i)}+\sigma_{t}\epsilon_{t,\mathcal{N}_{j}^{r}},t)+\sigma_{t}^{-1}\epsilon_{t,j}\right\|^{2}\right]\mathrm{d}t.$ (3.7)
@@ 3.2.1 Error Decomposition
$\mathbb{E}_{\mathsf{Q}}\left[\exp\left(\frac{1}{2}\int_{0}^{T-{\underline{t}}}\left\|\widehat{s}(Y_{t},T-t)-s(Y_{t},T-t)\right\|^{2}\mathrm{d}t\right)\right]<\infty,$ (3.10)
@@ 3.2.2 Localized Score Function
$s^{*}=\operatorname*{arg\,min}_{s_{\theta}\in\mathscr{H}_{r}^{*}}\int_{\underline{t}}^{T}\int\left\|s_{\theta}(x,t)-s(x,t)\right\|^{2}p_{t}(x)\mathrm{d}x\mathrm{d}t$ (3.12)
$\Leftrightarrow~\forall j,~s^{*}_{j}(x,t)=u_{j}^{*}(x_{\mathcal{N}_{j}^{r}},t),\quad u_{j}^{*}=\operatorname*{arg\,min}_{u_{\theta,j}\in\mathscr{U}_{j}}\int_{\underline{t}}^{T}\int\|u_{\theta,j}(x_{\mathcal{N}_{j}^{r}},t)-s_{j}(x,t)\|^{2}p_{t}(x)\mathrm{d}x\mathrm{d}t.$ (3.13)
@@
$|\mathcal{N}_{j}^{r}|=|\{i\in\mathbb{Z}^{\nu}:\left\|i\right\|_{1}\leq r\}|\leq(2r+1)^{\nu}<1+(3r)^{\nu}.$ (3.16)
@@
$C=2S\max\{1,m^{-1}\}\nu!\kappa^{2\nu+1}\log\kappa,\quad c=-2\log(1-\kappa^{-1}).$ (3.18)
@@
$\int_{0}^{T}\left\|s_{\theta}(x,t)-s(x,t)\right\|_{L^{2}(p_{t})}^{2}\mathrm{d}t\leq Cd(r+1)^{\nu}\mathrm{e}^{-c(r+1)},$ (3.20)
@@ 3.3.1 Equivalent to Diffusion Models for Marginals
$\begin{split}\mathcal{L}_{j}(u_{\theta,j})=~&\mathbb{E}_{x_{0,\mathcal{N}_{j}^{r}}\sim p_{0}}\int_{\underline{t}}^{T}\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|u_{\theta,j}(\alpha_{t}x_{0,\mathcal{N}_{j}^{r}}+\sigma_{t}\epsilon_{t,\mathcal{N}_{j}^{r}},t)+\sigma_{t}^{-1}\epsilon_{t,j}\right\|^{2}\right]\mathrm{d}t\\ =~&\mathbb{E}_{x_{0,\mathcal{N}_{j}^{r}}\sim p_{0}}\int_{\underline{t}}^{T}\mathbb{E}_{x_{t,\mathcal{N}_{j}^{r}}\sim p_{t|0}(x_{t,\mathcal{N}_{j}^{r}}|x_{0,\mathcal{N}_{j}^{r}})}\left[\left\|u_{\theta,j}(x_{t,\mathcal{N}_{j}^{r}},t)-\nabla_{j}\log p_{t|0}(x_{t,\mathcal{N}_{j}^{r}}|x_{0,\mathcal{N}_{j}^{r}})\right\|^{2}\right]\mathrm{d}t\\ =~&\int_{\underline{t}}^{T}\mathbb{E}_{x_{t,\mathcal{N}_{j}^{r}}\sim p_{t}}\left[\left\|u_{\theta,j}(x_{t,\mathcal{N}_{j}^{r}},t)-u_{j}^{*}(x_{t,\mathcal{N}_{j}^{r}},t)\right\|^{2}\right]\mathrm{d}t+{\rm const}.\end{split}$ (3.22)